# Crossroads of Continents: Automated Artifact Extraction for Cultural Adaptation with Large Multimodal Models

### v1

@@ C.1 Dataset details
car, cups, mugs and glasses, family snapshots, front door, home, kitchen, plate of food, social drink, wall decoration and wardrobe.
@@
car, cups, mugs and glasses, family snapshots, front door, home, kitchen, plate of food, social drink, wall decoration and wardrobe.
@@
  • "id": The language code for Indonesian, which is primarily spoken in Indonesia, corresponds to the South-eastern Asia sub-region.
  • "sw": The language code for Swahili, used in countries such as Tanzania, Kenya, and Rwanda, is mapped to the Eastern Africa sub-region.
  • "ta": The language code for Tamil, spoken in India and Sri Lanka, is associated with the Southern Asia sub-region.
  • "tr": The language code for Turkish, which is the official language of Turkey, falls under the Western Asia sub-region.
  • "zh": The language code for Chinese, predominantly spoken in China, is linked to the Eastern Asia sub-region.
@@
C.2 Prompt details
@@
C.3 Human Study - Interface and Annotators
@@
All annotators are demographically located in the USA and are between 25-40 years old. Roughly 40% identify as female, and the rest as male. In terms of an education background, 40% of the annotators are graduate students, whereas the rest includes working professionals from different backgrounds and also computer science faculty. In total, we have 14 annotators, recruited from different computer science labs at an university and also from a diverse set of social connections for this study. All the annotators have agreed to consent for using this data for research purposes. Our study qualifies for exemption from IRB as no PII is involved.
General Instructions for Human Study
We use https://labelstud.io Tkachenko et al. (2020–2022) to perform our human study. For each annotator, we create 2 tabs corresponding to the 2 studies, and ask them to solve them in numerical order to avoid potential biases from seeing true labels first from the second study. Time taken to complete the first study is usually 2 hours, and the time taken to complete the second study is typically 30 minutes. All annotators will be compensated for their time with a \$20 gift card upon completion of the task.
@@
C.4 Additional Results
$\Delta_{1}$ $\Delta_{2}$
Greece $\rightarrow$ China $-8.48$ $5.07$
Greece $\rightarrow$ India $-3.47$ $0.11$
Greece $\rightarrow$ USA $-6.13$ $5.81$
Table 3: Delta values for cultural adaptation from source country to target country. A successful cultural adaptation is indicated by $\Delta_{1}<0$ and $\Delta_{2}>0$.
@@ C.4 Additional Results
More examples of edits using CultureAdapt
@@
(a) The counts of answers obtained from our human study about the appropriateness of generated images shows that mostly people agree or are neutral, with only 2 disagreements overall.
(b) We measure accuracy for the cultural awareness task at the country, subregion and region level and find increasing performance in this order across our annotator pool.
Figure 20: (a) Appropriateness of generated images as perceived by human study participants. (b) Accuracy of cultural awareness at various geographical levels across our annotator pool for a subset of randomly sampled Dalle Street images.
@@
C.5 LLM hyperparameters
Generation settings
  • DALL-E 3 images were generated for vivid and natural settings for standard quality and size $1024$ x $1024$
  • GPT-4 and GPT-4V generations were obtained for temperature $=0.7$, top_p $=0.95$, no frequency or presence penalty, no stopping condition other than the maximum number of tokens to generate, max_tokens $=300$.
  • LLaVA generations were obtained for temperature $=1.0$ and top_p $=1.0$, no penalties, and max_tokens $=128$. The reason for using a slightly higher temperature and top_p is to have more consistent outputs. In our initial experiments, LLaVA did not perform as well in terms of following instructions at the same temperatue setting as GPT-4V.
  • For Stable Diffusion, we use stabilityai/stable-diffusion-2-inpainting from Hugging Face, and replace the autoencoder with stabilityai/sd-vae-ft-mse. We also use a DPMSolverMultistepScheduler for speeding up the generation process. We add "intricate details. 4k. high resolution. high quality." to the end of our prompt to get high quality images.
Computation budget
  • Total emissions for API based models are estimated to be 25 kgCO2 eq, of which 100 percents were directly offset by the cloud provider. Total emissions for our on-premise GPU usage is estimated to be less than 5 kgCO2 eq. Estimations were conducted using the MachineLearning Impact calculator Lacoste et al. (2019).

### v2

@@ A.1 Dataset details
car, family snapshots, front door, home, kitchen, plate of food, cups/mugs/glasses, social drink, wall decoration, and wardrobe.
@@
car, family snapshots, front door, home, kitchen, plate of food, cups/mugs/glasses, social drink, wall decoration, and wardrobe.
@@
  • ‘‘id’’: The language code for Indonesian, which is primarily spoken in Indonesia, corresponds to the South-eastern Asia sub-region.
  • ‘‘sw’’: The language code for Swahili, used in countries such as Tanzania, Kenya, and Rwanda, is mapped to the Eastern Africa sub-region.
  • ‘‘ta’’: The language code for Tamil, spoken in India and Sri Lanka, is associated with the Southern Asia sub-region.
  • ‘‘tr’’: The language code for Turkish, which is the official language of Turkey, falls under the Western Asia sub-region.
  • ‘‘zh’’: The language code for Chinese, predominantly spoken in China, is linked to the Eastern Asia sub-region.
@@
A.2 Prompt details
A.2.1 Dalle Street generation
Prompt for data generation
We use a simple template to prompt DALL-E 3 to generate images for a particular combination of country and category (Figure 14).
A.2.2 Cultural Awareness classifier
Prompt for classification of images
We use a simple prompt to generate names of subregions from models when provided an image.
Classification Prompt Strictly follow the United Nations geoscheme for subregions. Which geographical subregion of the United Nations geoscheme is this image from? Make an educated guess. Answer in one to three words.
Figure 15: We use a simple prompt to classify images in the data, by generating subregion labels.
A.2.3 Object Detection
@@
Figure 16: We use a detailed prompt for GPT-4V to extract objects, colors and counts from images generated with DALL-E 3.
@@
Figure 17: We use a detailed prompt for GPT-4 to process the dictionaries generated with the previous prompt into a simplified list along with some parsing rules to ensure correctness of the data structure.
A.3 LLM hyperparameters
A.3.1 Generation settings
  • DALL-E 3 images are generated for vivid and natural settings for standard quality and size $1024\times 1024$
  • GPT-4 and GPT-4V generations are obtained for temperature $=0.7$, top_p $=0.95$, no frequency or presence penalty, no stopping condition other than the maximum number of tokens to generate, max_tokens $=300$.
  • LLaVA generations are obtained for temperature $=1.0$ and top_p $=1.0$, no penalties, and max_tokens $=128$. The reason for using a slightly higher temperature and top_p is to have more consistent outputs. In our initial experiments, LLaVA did not perform as well in terms of following instructions at the same temperatue setting as GPT-4V.
  • For Stable Diffusion, we use stabilityai/stable-diffusion-2-inpainting from Hugging Face, and replace the autoencoder with stabilityai/sd-vae-ft-mse. We also use a DPMSolverMultistepScheduler for speeding up the generation process. We add ‘‘intricate details. 4k. high resolution. high quality.’’ to the end of our prompt to get high quality images.
A.3.2 Computation budget
  • Total emissions for API based models are estimated to be 25 kgCO2 eq, of which 100 percents were directly offset by the cloud provider. Total emissions for our on-premise GPU usage is estimated to be less than 10 kgCO2 eq. Estimations are conducted using the MachineLearning Impact calculator Lacoste et al. (2019).
A.4 Human Study - the Annotators
@@
All annotators have different demographic backgrounds (but are physically located in the USA currently) and are between 25-40 years old. Together, they are native to or have resided in more than countries and all 4 major regions from our dataset, and thus represent a strong sample of opinions. Roughly 40% identify as female, and the rest as male. In terms of an education background, 40% of the annotators are graduate students, whereas the rest includes working professionals from different backgrounds and also computer science faculty. In total, we have 14 annotators, recruited from different computer science labs at an university and also from a diverse set of social connections for this study. All the annotators have agreed to consent for using this data for research purposes. Our study qualifies for exemption from IRB as no PII is involved.
How many studies do we have
  1. 1.
    Human study to verify the quality of generated images
  2. 2.
    Human study to understand performance on our benchmark
  3. 3.
    Human validation for checking if artifacts are hallucinated
  4. 4.
    Human study to check if artifacts are culturally relevant
  5. 5.
    Human study for verifying people count associations
  6. 6.
    Human study for layout preservation after editing using CultureAdapt
  7. 7.
    Human study for cultural relevance after editing using CultureAdapt
Total number of annotations
To validate our dataset and human baseline on cultural awareness, we have $14$ people, each annotating $300$ samples, resulting in $4200$ annotations for one study and $8400$ in total. For each of the other three studies (excluding the studies for edits), we have three people annotating $100$ samples each for a study, resulting in $900$ annotations in total. Finally, for the two editing-related studies, we have three people each looking at three images per sample (source image, edited image using our method, edited image using another method) for $100$ samples, resulting in $900$ annotations per study or $1800$ in total. Combining all this, we have $8200+900+1800=10900$ annotation items to back the comprehensive quantitative evaluations we perform for each study in our paper.
A.5 Human Study - the Interfaces
General Instructions for Study 1, 2
We use https://labelstud.io Tkachenko et al. (2020–2022) to perform our human study. For each annotator, we create 2 tabs corresponding to the 2 studies, and ask them to solve them in numerical order to avoid getting influenced from seeing true labels first from the second study. Time taken to complete the first study is usually 2 hours, and the time taken to complete the second study is typically 30 minutes. All annotators will be compensated for their time with a \$20 gift card upon completion of the task.
A.5.1 Study 1: verify the quality of generated images
@@
A.5.2 Study 2: human performance on our cultural awareness benchmark
@@
A.5.3 Study 3, 4: artifact hallucination and cultural relevance
We provide an interface (Figure 20) that includes a multi-choice correct question answering setting, where we first ask a question about the artifacts that are present in the image and then ask if those artifacts are culturally relevant. For selecting these options, we sample randomly from all artifacts extracted for the accompanying image. For this study, we also provide the name of the concept class that the image is supposed to represent and the country, so that annotators have an idea about the object categories that they might be looking for. To account for cases where annotators might not be sure about any object and still choose one, in our analysis we only consider those cases where atleast two artifacts are marked as present. Responses to the second question is somewhat subjective as it depends on the cultural backgrounds of the annotators, but as we see in our analysis, the responses mostly agree and find more than half of the artifacts to be relevant.
Figure 20: Annotation Interface for Study 3,4
A.5.4 Study 5: people count associations
We used an identical interface to the one for Study 3,4 to carry out this study, but instead of providing multiple correct options, we provide a direct True/False setting for annotators to answer if the count of people shown in the given image is approximately correct, i.e. falls in the correct count bucket (1-5, 5-10, more than 10). Our analysis reveals that people mostly agree with the count buckets generated, but on corroborating patterns with actual population statistics, the ordering of countries look slightly different, so we do not make any correlations about that in our analysis.
A.5.5 Study 6, 7: layout preservation and cultural relevance after editing
We use an interface (Figure 21) that shows the name of the source and target countries, the source image, and two edited images - one from our method CultureAdapt and one from cap-edit. We ask a somewhat subjective question regarding which image the annotator prefers, but responses vary widely for this question and do not correlate well across annotators, so we do not use these results for further analysis. For each of the edited images, we ask three questions - (1) if the edit maintains structural layout, (2) if the edit makes the image more culturally relevant to the target country and (3) if the edited image might be toxic for people from target country. All of these questions require a response on a Likert scale, which we design using a system of stars, where users can select anywhere between 1 and 5 stars. Less than 1% of images are marked toxic across all annotations and annotators agree strongly on this metric. We report the analysis for the other two metrics in the main paper.
Figure 21: Annotation Interface for Study 6,7
A.6 Additional Results
A.6.1 Task 1 - Cultural awareness
Income distribution performance for LLaVA
We show the performance of LLaVA on the income quartile distribution basis for Dollar Street.
Confusion matrix set 1
We show confusion matrices for both LLaVA and GPT-4V on Dalle Street images.
Confusion matrix set 2
We show confusion matrices for both LLaVA and GPT-4V on Dollar Street images.
Confusion matrix set 3
We show confusion matrices for both LLaVA and GPT-4V on MaRVL images.
A.6.2 Task 2 - Artifacts
@@
Table 4: Salient Artifact Statistics (adj indicates descriptors like color, etc are part of the artifact name, whereas no_adj indicates the artifact name does not have such descriptors)
Color associations on DALL-E 3 images
We find that countries are more likely to be associated with particular colors, with some showing prominently strong associations.
@@
People-count associations on DALL-E 3 images
Distributions validated by humans do not always correlate with actual population statistics.
@@
Interesting associations
We show examples of interesting associations identified by models and humans at the country level, for our artifact extraction task.
@@
A.6.3 Task 3 - Edits
More examples of edits using CultureAdapt
CultureAdapt quantitative metrics
Source-Target Pair Similarity SSIM Metric $M_{1}$ Metric $M_{2}$
cap-edit CultureAdapt cap-edit CultureAdapt cap-edit CultureAdapt
Brazil-India 0.96 0.93 0.72 0.68 0.95 0.95
Brazil-Nigeria 0.96 0.93 0.62 0.47 0.89 0.91
Brazil-Turkey 0.96 0.93 0.69 0.52 0.95 0.94
Brazil-USA 0.96 0.93 0.28 0.35 0.91 0.83
Average 0.96 0.93 0.58 0.51 0.93 0.91
India-Brazil 0.97 0.94 0.63 0.60 0.96 0.90
India-Nigeria 0.96 0.94 0.69 0.68 0.94 0.90
India-Turkey 0.97 0.94 0.56 0.57 0.92 0.89
India-USA 0.96 0.94 0.67 0.63 0.93 0.88
Average 0.97 0.94 0.63 0.62 0.94 0.89
Nigeria-Brazil 0.96 0.92 0.29 0.41 0.76 0.85
Nigeria-India 0.96 0.92 0.62 0.67 0.87 0.93
Nigeria-Turkey 0.96 0.91 0.66 0.63 0.91 0.93
Nigeria-USA 0.96 0.92 0.50 0.60 0.90 0.91
Average 0.96 0.92 0.51 0.58 0.86 0.90
Turkey-Brazil 0.97 0.94 0.46 0.41 0.89 0.84
Turkey-India 0.97 0.95 0.62 0.59 0.88 0.88
Turkey-Nigeria 0.97 0.94 0.67 0.64 0.93 0.90
Turkey-USA 0.96 0.94 0.29 0.43 0.88 0.89
Average 0.97 0.94 0.51 0.51 0.89 0.88
USA-Brazil 0.97 0.94 0.53 0.46 0.88 0.89
USA-India 0.98 0.94 0.18 0.62 0.58 0.91
USA-Nigeria 0.98 0.94 0.20 0.46 0.61 0.84
USA-Turkey 0.98 0.94 0.21 0.46 0.64 0.87
Average 0.97 0.94 0.28 0.50 0.68 0.88
Overall Average 0.97 0.94 0.50 0.54 0.85 0.89
Table 7: Mean Similarity (SSIM) Scores, Metric $M_{1}$, and Metric $M_{2}$ for cap-edit and CultureAdapt grouped by Source and Target Country